# Machine Learning in Fundamental Physics Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature

### v1

@@
Abstract
We propose a unified framework that brings together statistical physics, coding theory, and algebraic topology for multi-class image classification. First, each input image is passed through a frozen MobileNetV2 backbone to obtain a high-dimensional feature vector $\mathbf{h}\in\mathbb{R}^{1280}$. The components of $\mathbf{h}$ are interpreted as Ising spins placed on the vertices of a sparse Multi-Edge Type quasi-cyclic LDPC (MET-QC-LDPC) graph. The resulting Random-Bond Ising Model (RBIM) is operated at its Nishimori temperature $\beta_{N}$, which we identify as the unique point where the smallest eigenvalue of the Bethe–Hessian matrix $H_{\beta,J}$ vanishes: $\lambda_{\min}\!\bigl{(}H_{\beta_{N},J}\bigr{)}=0.$ At this critical temperature class separability is maximized. Our first theoretical contribution is an exact correspondence between local trapping sets in the Tanner graph of a QC-LDPC code and topological invariants of the underlying feature manifold. We prove that each elementary cycle generates a pole of the Ihara–Bass zeta function, which appears as an isolated eigenvalue of both the non-backtracking operator and $H_{\beta,J}$. These spectral signatures encode $\mathbb{Z}/2$-torsion (Betti numbers, bordism classes) that act as topological defects in the embedding. The second contribution is a practical algorithm for estimating $\beta_{N}$. By evaluating $\lambda_{\min}(H_{\beta,J})$ at three trial temperatures, fitting a quadratic interpolant, and applying one Newton correction, the method converges in roughly nine Arnoldi iterations—approximately a six-fold speed-up over standard bisection. Guided by the topological analysis we design two MET-QC-LDPC graph ensembles (a spherical family with a single circulant ring and a toroidal family with two independent rings). Permanent and Bethe–permanent bounds are used to suppress harmful trapping sets during protograph and parity-check matrix optimization. The resulting spectral embeddings compress the original 1280-dimensional MobileNetV2 features to $32$ dimensions for a 10-class ImageNet subset (ImageNet-10), and $64$ dimensions for a 100-class ImageNet subset (ImageNet-100). Despite this compression we achieve 98.7% top-1 accuracy on ImageNet-10 (spherical graphs) and 82.7% top-1 accuracy on ImageNet-100 using a three-graph ensemble, while reducing the number of parameters and memory footprint by factors of $40$ and $20$, respectively. In summary, controlling graph design through topological invariants yields highly efficient, physics-inspired embeddings that retain state-of-the-art classification performance.
@@ 1 Introduction
Graph-based spectral methods have recently shown great promise for machine learning tasks. In particular, modeling feature vectors as spins (nodes) on a sparse graph under a Gibbs distribution (a Random-Bond Ising Model) enables powerful clustering and classification. For example, Nishimori and Bethe free-energy ideas have been used to tune graph-based classifiers ([1] , and prior work demonstrated that sparse QC-LDPC graphs with Nishimori-weighted edges can significantly improve clustering of high-dimensional CNN features, [2]. However, most of these successes have been in synthetic or binary settings (e.g. distinguishing GAN-generated images), while natural multi-class image datasets pose additional challenges: the feature manifolds can have nontrivial topology and curvature that corrupt spectral embeddings, and standard methods struggle to capture nonlinear class boundaries. In this work, we develop a rigorous, end-to-end framework that unifies ideas from statistical physics, coding theory, and topology to tackle these challenges. Our contributions are threefold. We derive explicit relationships between graph-theoretic trapping sets (small subgraphs in Tanner graphs) and spectral/topological invariants. In particular, we link closed cycles in the graph to zeros and poles of the Ihara–Bass zeta function, and show how these induce isolated eigenvalues in the Bethe–Hessian and non-backtracking operators. This analysis reveals that trapping sets act as torsion defects in the feature manifold, which can be characterized by topological invariants (Betti numbers, quadratic-form signature, etc.) and by $K$-theory indices. We introduce a novel algorithm to estimate the Nishimori temperature in graph-based RBIMs. By fitting a low-degree Lagrange interpolant to the smallest Bethe–Hessian eigenvalue and imposing a stability constraint, our method converges reliably to the critical $\beta_{N}$ even under complex geometric constraints. This yields a provably optimal spin-glass/paramagnetic threshold for clustering. We apply ensembles of three expert quasi-cyclic LDPC graphs to real image classification benchmarks (ImageNet-10 and ImageNet-100). Starting from 1280-dim MobileNetV2 features, we use these structured sparse graphs to perform a spectral embedding that compresses features to 32 (resp. 64) dimensions per class. Using our Nishimori-tuned RBIM weights and ensemble majority voting, we achieve 98.7% accuracy on ImageNet-10 and 82.7% on ImageNet-100 (comparable to a 64-D dense embedding), while reducing parameters by a factor of 20×. The remainder of the paper is organised as follows. Section 2 recalls the RBIM and the Nishimori condition. Section 3 introduces quasi-cyclic LDPC graphs and their multi-edge extensions. Topological consequences of trapping sets are discussed in Section 4. The temperature estimator is presented in Section 4.6. Experimental results on ImageNet are reported in Section 5, and we conclude with future directions in Sections  6,  Conclusion.
@@ 2 Random Bond Ising Models
Let $\mathcal{G}=(V,E)$ be an undirected graph with $|V|=n$. Assign a binary spin variable $s_{i}\in\{-1,+1\}$ to each vertex $i\in V$ and denote the spin configuration by $\mathbf{s}=(s_{1},\dots,s_{n})^{\!\top}$. For a symmetric coupling matrix $J=(J_{ij})_{i,j=1}^{n}$, the Hamiltonian of the Random-Bond Ising Model is
@@ 2.1 Nishimori condition and phase transition
the Nishimori condition holds. At $\beta=\beta_{N}$ the model exhibits a paramagnetic–spin-glass transition, and many observables (e.g. magnetisation, overlap) become analytically tractable.
@@
Thus estimating $\beta_{N}$ reduces to solving a one-dimensional root–finding problem.
@@ 2.2 Bethe free energy and the Bethe–Hessian
Figure 1: In (a) and (b) we see cluster structures (community) from phase transaction described by adjacency matrices and network related graph, Fig. 20 from [5]
3 Multi-Edge Type Quasi-Cyclic LDPC Graphs
Low-density parity-check (LDPC) codes naturally define sparse bipartite graphs. A Quasi-Cyclic LDPC (QC-LDPC) codes provide a structured and hardware-friendly subclass defined by a quasi-cyclic parity-check matrix $H$ [8]. An $(N,K)$ QC-LDPC code consists of $N$ total codeword bits, with $K$ information bits and NKN-K parity bits. The associated Tanner graph is described by a parity-check matrix $H\in\mathbb{F}_{2}^{mL\times nL}$, constructed from square blocks of size $L\times L$, where each block is either a zero matrix or a circulant permutation matrix (CPM) [9, 10].
@@ A CPM $P\in\{0,1\}^{L\times L}$ is defined as:
Figure 2: (Left) Tanner graph corresponding to the parity-check matrix $H_{1}$, composed by 2 and 3 ring of size 7, Fig.2 [11]. (Right) Multi-graph protograph $M(H_{2})$ representation of the QC-LDPC code $H_{2}$.
@@
4 Topological Signatures of Trapping Sets
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{10}$$c_{11}$$c_{12}$
Figure 3: Graphical representation of TS(4, 2) (left), TS(4, 6) (center) and TS(9, 2) (right)
4.1 Spectral manifestation of trapping sets
[…]
4.2 Ihara–Bass zeta function
For a finite graph $G$ let $\mathcal{B}$ denote the non-backtracking operator. The Ihara–Bass zeta function is defined by $\zeta_{G}(u)=\prod_{[C]}\bigl{(}1-u^{\,\ell(C)}\bigr{)}^{-1},$ where the product runs over equivalence classes $[C]$ of primitive cycles and $\ell(C)$ is the length of a cycle,  [19] . Poles of $\zeta_{G}$ occur at reciprocals of eigenvalues of $\mathcal{B}$; therefore each closed cycle contributes a factor $(1-u^{\ell})^{-1}$. When $G$ contains a trapping set, the associated cycles generate poles in $\zeta_{G}(u)$, which manifest as isolated eigenvalues of both $\mathcal{B}$ and the Bethe–Hessian $H_{\beta,J}$. These poles can be interpreted as $\mathbb{Z}/2$-torsion elements: they correspond to non-trivial $1$-cycles ($\beta_{1}>0$) in the underlying feature manifold. Removing trapping sets eliminates the corresponding poles and reduces the number of low-energy modes.
4.3 Bordism triviality and topological obstructions
Let $G=(V,E)$ be a simple graph and let $\phi:V\rightarrow\mathbb{R}^{d}$ embed its vertices as points of a smooth manifold $M_{G}:=\phi(V)\subset\mathbb{R}^{d}$, [20]. We call $G$ bordism-trivial if $M_{G}$ is null–bordant in the appropriate bordism group, i.e. there exists a compact $(d+1)$-manifold $W$ with $\partial W=M_{G}\qquad(\text{or }\partial W=M_{G}\sqcup(-M_{G})\text{ for the unoriented case}),$ and all characteristic numbers of $M_{G}$ vanish: $w_{k}(M_{G})=0,\;k>0,\qquad p_{j}(M_{G})=0,\;j\geq 1.$ Equivalently $M_{G}]=0$ in the oriented bordism group $\Omega^{SO}_{d}$ (and in $\Omega^{O}_{d}$ when orientation is ignored). For a multi-class problem let $\{M_{i}\}_{i=1}^{K}$ be the manifolds obtained from the feature vectors of each class. A bordism obstruction between classes $i$ and $j$ occurs if the disjoint union $M_{ij}:=M_{i}\sqcup(-M_{j})$ is not null-bordant in $\Omega^{SO}_{d}$ (or $\Omega^{O}_{d}$). A convenient detection criterion is the non-vanishing of a characteristic number, e.g. $\exists\,k>0:\;w_{k}[M_{ij}]\neq 0\quad\text{or}\quad\exists\,j:\;p_{j}[M_{ij}]\neq 0.$ If such an obstruction exists for any pair of classes, no graph-based embedding can achieve perfect separation in the high-dimensional feature space, [21].
4.4 Higher signatures and KK–theory
Given a cell complex $\mathcal{M}_{G}$ derived from the graph, let $L_{k}=d_{k-1}^{*}d_{k-1}+d_{k}d_{k}^{*}$ be the combinatorial Hodge Laplacian acting on $k$-cochains, [22]. The kernel dimension satisfies $\dim\ker L_{k}=\beta_{k},$ so that spectral information directly yields Betti numbers.
For any cohomology class $x\in H^{*}(B\pi_{1}(\mathcal{M}_{G});\mathbb{Q})$, the higher signature $\sigma_{x}(\mathcal{M}_{G})=\bigl{\langle}L(\mathcal{M}_{G})\smile x,\,[\mathcal{M}_{G}]\bigr{\rangle}$ pairs the Hirzebruch $L$-class with the fundamental class,  [23]. By the Novikov conjecture, $\sigma_{x}\!\!\pmod{2}$ detects non-trivial $\mathbb{Z}/2$-torsion, hence odd-length trapping cycles render some higher signature non-zero.
The Dirac operator on the bipartite graph, $\mathcal{D}=\begin{pmatrix}0&A_{\!vn}\\ A_{\!vn}^{\!\top}&0\end{pmatrix}$, has spectrum symmetric about zero, [24, 25]. Its analytic index equals the difference of dimensions of positive and negative eigenspaces, which, by the Atiyah–Singer index theorem, coincides with $\beta_{0}-\beta_{1}$, [26]. Thus a non-zero index signals the presence of topological defects (e.g. odd cycles).
In $K$-theory, the Kasparov matrix $D_{\mathrm{Kas}}=\begin{pmatrix}0&S&T\\ S^{\!\top}&0&0\\ T^{\!\top}&0&0\end{pmatrix}$, built from suitable incidence matrices $S$ and $T$, satisfies: $K_{0}=\dim\ker D_{\mathrm{Kas}}^{2}$ (counts zero-modes, i.e. $\beta_{0}$); $K_{1}=\operatorname{rank}D_{\mathrm{Kas}}\bmod 2$ (detects $\mathbb{Z}/2$ torsion, equivalent to the Kervaire invariant), [27] . Non-trivial $K$-groups therefore provide a concise algebraic signature of the topological obstruction introduced by a trapping set, [28].
The continuous genus is a scalar number that captures how “twisted’’ or “curved’’ a graph $G$ (or, more generally, a network built from points and edges) is, by looking at the spectrum of a matrix derived from the graph $\widehat{A}(H)=\frac{1}{2\sqrt{n_{V}}}\Biggl{(}\sum_{\lambda_{i}\in\Lambda^{+}}\!\sqrt{\lambda_{i}}-\sum_{\lambda_{j}\in\Lambda^{-}}\!\sqrt{-\lambda_{j}}\Biggr{)}$, where $n_{V}$ are number of vertices in the graph, [29]. The experimental study of how topological invariants influence spectral embedding is presented in Sec. 5.1.
4.5 Permanent and Bethe–permanent bounds
The permanent of an $m\times m$ matrix $\mathbf{B}=[b_{j,i}]_{j,i}$ over a commutative ring is $\operatorname{perm}(\mathbf{B})=\sum_{\sigma\in S_{m}}\;\prod_{j=1}^{m}b_{\,j,\sigma(j)},$ i.e. the determinant without the sign factor $\operatorname{sgn}(\sigma)$. Exact evaluation of (4.5) is #P-complete, consequently a number of approximation and fast-evaluation schemes have been proposed [30]. The minimum Hamming distance $d_{\min}$ of a QC-LDPC code can be bounded by permanents of its weight matrix [31]. For a parity-check matrix $H(x)$ let $\mathbf{A}=\operatorname{wt}(H(x))$ denote the entrywise Hamming weight. The code distance upper bound, $TS(a=d_{\min},0)$:
where $\operatorname{min}^{*}$ denotes the minimum over all $(v+1)$-subsets. Replacing the permanent by its Bethe approximation yields an upper bound on the minimum pseudoweight, which controls the harmfulness of $TS(a,b>0)$,  [32]. The Bethe-permanent of a non-negative matrix $\mathbf{B}$ is defined as $\operatorname{perm}_{\!B}(\mathbf{B})=\exp\!\Bigl{(}-\min_{q\in\mathcal{Q}}\;F_{\rm Bethe}(q)\Bigr{)}$, where $F_{\rm Bethe}$ is the Bethe free energy and $\mathcal{Q}$ denotes the set of factorised marginals, [33]. In our construction we maximize weight for both bounds during protograph and parity-check optimization using EMD/ACE maximization, (Alg. 2, [2]), thereby suppressing low-weight TS and improving the Bethe–Hessian spectrum, Eq. 5.
@@ 4.5 Permanent and Bethe–permanent bounds
Figure 4: Dependence of the smallest eigenvalue $\lambda_{\min}$ on the temperature parameter $\beta$. The red curve shows the polynomial part, the blue curve the tail, and the black dashed line is a quadratic approximation to the polynomial part. The coefficient of determination for the fit is $R^{2}=0.9998$.

### v2

@@
Machine Learning in Fundamental Physics
Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature
V.S. Usatyuk L@Lcrypto.com Research and Development Department, T8 LLC, Moscow, 107076 Russia Department of Computer Engineering, South-West State University, Kursk, 305040 Russia    D.A. Sapozhnikov Research and Development Department, T8 LLC, Moscow, 107076 Russia    S.I. Egorov Department of Computer Engineering, South-West State University, Kursk, 305040 Russia
($x$.xx.2025)
Аннотация
Modern multi‑class image classification relies on high‑dimensional convolutional‐neural‑network (CNN) feature vectors, which incur large memory and computational costs and provide little insight into the geometry and properties of the underlying vision data manifold. Existing graph‑based spectral classifiers have shown promise on synthetic or binary classification problems, but they degrade on natural images with multi-classes because the feature manifolds possess non‑trivial topology that destroys class separability when conventional graphs are used. We introduce a physics‑inspired pipeline in which frozen MobileNetV2 features are interpreted as Ising spins placed on the vertices of a sparse Multi‑Edge Type quasi‑cyclic LDPC (MET‑QC‑LDPC) graph, thereby defining a Random‑Bond Ising Model (RBIM). The model is operated at its Nishimori temperature—identified as the unique point where the smallest eigenvalue of the Bethe–Hessian matrix vanishes. Two methodological pillars underpin the design. An exact spectral–topological correspondence that links local trapping sets in the Tanner graph to topological invariants (via poles of the Ihara–Bass zeta function), enabling systematic suppression of harmful substructures through permanent and Bethe‑permanent bounds. Such a harmful subgraph reduces top‑1 performance by more than a factor of four in multi‑class classification. A fast quadratic–Newton estimator for the Nishimori temperature that converges in roughly nine Arnoldi iterations, providing a sixfold speed‑up over standard bisection. This method enables spectral graph embedding training on large‑scale datasets such as ImageNet‑100. The resulting graph ensembles compress the original 12801280‑dimensional MobileNetV2 representation to 32 dimensions for a ten‑class ImageNet subset (ImageNet‑10) and to 64 dimensions for a hundred‑class subset (ImageNet‑100). Despite this compression we achieve $98.7\,\%$ top‑1 accuracy on ImageNet‑10, and up to 84.92% top‑1 accuracy on ImageNet‑100 using three‑graph soft ensemble. Compared with MobileNetV2, our hard‑ensemble approach yields a 0.10% increase in top‑1 accuracy while reducing the number of FLOPs by a factor of 2.67. In contrast, when benchmark against ResNet‑50, the soft‑ensemble variant suffers only a 1.09% drop in top‑1 accuracy yet cuts FLOPs by 29-times, thereby achieving a favorable accuracy–efficiency trade‑off. The novelty of this work lies in (a) establishing a rigorous link between graph trapping sets and algebraic‑topological defects, (b) providing an efficient Nishimori‑temperature estimator, and (c) demonstrating that topology‑guided LDPC graph embedding yields highly compressed classifiers.
@@ 1 Introduction
Deploying deep neural networks on bandwidth‑constrained, latency‑sensitive platforms (mobile phones, autonomous vehicles, remote sensors) requires that the classifier head – the linear layer mapping backbone embeddings to class scores – be both memory‑light and communication‑efficient. When the number of categories $C$ grows to hundreds or thousands (e.g., ImageNet‑100/1000), the head’s parameter count ($d\times C$) dominates overall resource consumption. A lots of papers has attempted to alleviate this bottleneck, reporting the following empirical outcomes: Linear dimensionality reduction (e.g., Probabilistic PCA) can shrink the activation dimension before classification, but at the cost of a substantial drop in top‑1 accuracy—up to 41.33% relative loss [1, 2]. Low‑rank factorizations of the weight matrix (singular‑value truncation) have yielded modest compression on simple benchmarks such as CIFAR‑10 and STL‑10, yet they do not translate into competitive performance on larger datasets [3]. Deep auto‑encoders that learn a non‑linear manifold for the classifier weights achieve higher compression ratios, but their training demands extensive label‑rich data, and the reconstruction objective does not directly optimise classification accuracy; over‑compression often erodes fine‑grained class information [3]. Cross‑Stage‑Partial (CSP) architectures reduce the backbone embedding size from the typical $d=1280$ (MobileNetV3) to $d=256$, thereby cutting classifier parameters by roughly 88% for ImageNet‑100 and raising top‑1 accuracy from 81.26% to 82.64% after dropout regularization [5]. Despite this improvement, CSP‑based models still lag behind larger CNNs (ResNet‑50: 86.69%) and Vision Transformers ($\approx 86.01\%$) on the same benchmark and lack a flexible mechanism for head ensembling that could trade latency for higher accuracy. Collectively, these approaches either sacrifice predictive performance or impose computational demands incompatible with real‑time, on‑device deployment, leaving an unmet need for a compact yet accurate classifier head.
We address this gap by re‑conceptualizing the classification head as a statistical‑physics system. Each class weight vector is treated as a spin on a sparse graph whose edges carry random couplings, forming a Random‑Bond Ising Model (RBIM). By enforcing the Nishimori condition and exploiting Bethe free‑energy approximations, we tune edge weights so that the resulting spin ensemble undergoes a sharp paramagnetic–ferromagnetic transition precisely at the decision boundary. Building on recent findings that quasi‑cyclic low‑density parity‑check (QC‑LDPC) graphs with Nishimori‑weighted edges improve binary clustering of CNN descriptors [7], we extend the construction to multi‑class settings.
The key outcomes of this strategy are: Extreme compression with competitive accuracy. The $1\,280\times C$ classifier matrix is compressed to a 32‑dimensional (for $C=10$) or 64‑dimensional (for $C=100$) spectral embedding per class. On ImageNet‑10 the method attains 98.7% top‑1 accuracy, and on ImageNet‑100 it reaches 82.7%; a soft ensemble of three QC‑LDPC graphs further lifts performance to 84.9% top‑1, all while requiring only matrix–vector products at inference time. Spectral–topological correspondence for trapping sets: We prove that small subgraphs (trapping sets) in the Tanner graph of a QC‑LDPC code generate isolated eigenvalues in both the Bethe‑Hessian and non‑backtracking operators, manifesting as torsion defects on the classifier manifold. These defects are quantifiable via classical topological invariants (Betti numbers) and $K$-theoretic indices. Robust Nishimori temperature estimator: A low‑degree Lagrange interpolant fitted to the smallest eigenvalue of the Bethe‑Hessian yields a monotonic, stable estimate of the critical Nishimori inverse temperature $\beta_{N}$. The algorithm converges reliably even under severe geometric distortion, providing an optimal spin‑glass/paramagnetic threshold for image clustering (unsupervised classification). Scalable validation: Experiments on ImageNet-10 and ImageNet-100 datasets confirm that the proposed graph‑based embedding satisfies real‑time constraints without sacrificing accuracy, outperforming prior compression schemes. In summary, our work introduces a fundamentally different paradigm—leveraging sparse spin‑glass models and spectral graph theory—to compress deep classification heads. This yields a highly compact, accurate, and inference‑efficient solution that bridges the performance–resource gap left by existing methods. The remainder of the paper is organised as follows: Section 2 reviews the RBIM and Nishimori condition; Section 3 details QC‑LDPC graph constructions; Appendix A analyses trapping‑set topology; Section 3.1 presents the temperature‑estimation algorithm; Section 4 reports experimental results. Sections 5 and Conclusion discuss future directions and conclude.
@@ 2 Random Bond Ising Models
Let $\mathcal{G}=(V,E)$ be an undirected graph with $|V|=n$. Assign a binary spin variable $s_{i}\in\{-1,+1\}$ to each vertex $i\in V$ and denote the spin configuration by $\mathbf{s}=(s_{1},\dots,s_{n})^{\!\top}$. For a symmetric coupling matrix $J=(J_{ij})_{i,j=1}^{n}$, the Hamiltonian of the Random‑Bond Ising Model is
@@ 2.1 Nishimori condition and phase transition
the Nishimori condition holds. At $\beta=\beta_{N}$ the model exhibits a paramagnetic–spin‑glass transition, and many observables (e.g. magnetisation, overlap) become analytically tractable.
@@
Thus estimating $\beta_{N}$ reduces to solving a one‑dimensional root–finding problem.
@@ 2.2 Bethe free energy and the Bethe–Hessian
Рис. 1: In (a) and (b) we see cluster structures (community) from phase transaction described by adjacency matrices and network related graph, Fig. 20 from [10]
3 Multi-Edge Type Quasi‑Cyclic LDPC Graphs
Low‑density parity‑check (LDPC) codes naturally define sparse bipartite graphs. A Quasi-Cyclic LDPC (QC-LDPC) codes provide a structured and hardware-friendly subclass defined by a quasi-cyclic parity-check matrix $H$ [13]. An $(N,K)$ QC-LDPC code consists of $N$ total codeword bits, with $K$ information bits and NKN-K parity bits. The associated Tanner graph is described by a parity-check matrix $H\in\mathbb{F}_{2}^{mL\times nL}$, constructed from square blocks of size $L\times L$, where each block is either a zero matrix or a circulant permutation matrix (CPM) [14, 15].
@@ A CPM $P\in\{0,1\}^{L\times L}$ is defined as:
Рис. 2: (Left) Tanner graph corresponding to the parity-check matrix $H_{1}$, composed by 2 and 3 ring of size 7, Fig.2 [16]. (Right) Multi-graph protograph $M(H_{2})$ representation of the QC-LDPC code $H_{2}$.
@@
Рис. 3: Dependence of the smallest eigenvalue $\lambda_{\min}$ on the temperature parameter $\beta$. The red curve shows the polynomial part, the blue curve the tail, and the black dashed line is a quadratic approximation to the polynomial part. The coefficient of determination for the fit is $R^{2}=0.9998$.